OpenStudy (anonymous):

Find the derivative of the function. y = ln(|csc(3x)|)

OpenStudy (anonymous):

You have to use chain rule to solve this.

OpenStudy (anonymous):

\[\frac{d}{dx}\ln(f(x))=\frac{f'(x)}{f(x)}\]

OpenStudy (anonymous):

-3 / (csc(3x)) * csc(3x)cot(3x)...I'm I right satellite?

OpenStudy (anonymous):

\[f(x)=\csc(3x),f'(x)=-3\csc(3x)\cot(3x)\]

OpenStudy (anonymous):

@nayadddo f' is in the numerator and f is in the denominator

OpenStudy (anonymous):

oh maybe you are right when you are done. you are faster than i am hold on

OpenStudy (anonymous):

Yeah that's what I meant what math type do you use to answer questions..is it the equation that they offer on this site?

OpenStudy (dumbcow):

should end up with -3cot(3x) i believe

OpenStudy (anonymous):

\[\frac{-3\csc(3x)\cot(3x)}{\csc(3x)}\] is a start

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

but i hate all these reciprocal trig functions so if you write everything out in terms of sine and cosine it will be much nicer.

OpenStudy (anonymous):

evidently the cosecants cancel, leaving you with \[-3\cot(3x)\] maybe best to leave it at that

OpenStudy (anonymous):

-3cot(3x)?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Satellite I have a question when x ix 0 it seems like the derivative is undefined for both csc(3x) and cot(3x) so can we still divide csc(3x) / csc(3x)...just wondering..